I am stuck with this problem. Please help.
The length of one of the equal legs of an isosceles triangle is 8cm less than 4 times the length of the base. If the perimeter is 29cm, find the lengths of the equal legs.

Answer by checkley71(8403)

2(4X-8)+X=29
8X-16+X=29
9X=29+16
9X=45
X=45/9
X=5 ANSWER FOR THE BASE.
4*5-8=20-8=12 FOR THE EQUAL SIDES.
PROOF
2(4*5-8)+5=29
2(20-8)+5=29
2*12+5=29
24+5=29
29=29